OpenStudy (anonymous):

y =2x + 6 4x - 2y = 5

OpenStudy (anonymous):

their is some mistake in the question...

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

ans comes to be zero..:p

OpenStudy (anonymous):

put value of y in equation 2 se what u got?

OpenStudy (anonymous):

substitute the first equation into the second: \[\large 4x-2y=5\Rightarrow 4x-2(2x+6)=5\Rightarrow 4x-4x-12=5\Rightarrow -12=5 \] this result of -12=5 is never true which means the solution to this system is NO solution. you can also determine this by solving for "y" in the second equation and you'll see the lines are parallel which means they will not intersect.... therefore no solution.